# The Hallucinations Leaderboard – An Open Effort to Measure Hallucinations in Large Language Models

### v1

@@ Abstract
Giwon Hong1thanks: Equal contribution. GH conducted the first draft of the paper and the analyses in Section 3. APG contributed to the first version of the leaderboard and corresponding blog post. RS contributed to the summarisation tasks. XD contributed to the knowledge memorisation tasks. LPB contributed to the writing and experimental design. MR contributed to prompt robustness evaluation and writing. PM created the first version of the leaderboard and corresponding blog post.  Aryo Pradipta Gema [^1] footnotemark: 1  Rohit Saxena  [^1] footnotemark: 1  Xiaotang Du [^1] footnotemark: 1  Ping Nie411footnotemark: 1 Yu Zhao [^1] footnotemark: 1  Laura Perez-Beltrachini1  Max Ryabinin3  Xuanli He2  Pasquale Minervini [^1] footnotemark: 1thanks: Corresponding authors. 1School of Informatics, University of Edinburgh 2Department of Computer Science, University College London  3Together AI 4School of Electronics Engineering and Computer Science, Peking University {first.last,lperez,p.minervini}@ed.ac.uk mryabinin0@gmail.com  xuanli.he@ucl.ac.uk   ping.nie@pku.edu.cn
@@ Acknowledgements
Experiments are being conducted mainly at the Edinburgh International Data Facility (EIDF) and on the internal clusters of the School of Informatics, University of Edinburgh. APG was supported by the United Kingdom Research and Innovation (grant EP/S02431X/1), UKRI Centre for Doctoral Training in Biomedical AI at the University of Edinburgh, School of Informatics. PM was partially funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1) and a donation from Accenture LLP; and is grateful to NVIDIA for the GPU donations. XH and PM are funded by an industry grant from Cisco. GW was supported by ILCC program (School of Informatics Funding Package) at the University of Edinburgh, School of Informatics. RS, XD, and YZ are supported in part by the UKRI Centre for Doctoral Training in Natural Language Processing, funded by UK Research and Innovation (grant EP/S022481/1) and the School of Informatics.

### v2

@@ Abstract
Giwon Hong1thanks: Equal contribution. GH conducted the first draft of the paper and the analyses in Section 3. APG contributed to the first version of the leaderboard and corresponding blog post. RS contributed to the summarisation tasks. XD contributed to the knowledge memorisation tasks. LPB contributed to the writing and experimental design. MR contributed to prompt robustness evaluation and writing. PM and CF created the first version of the leaderboard and corresponding blog post.  Aryo Pradipta Gema [^1] footnotemark: 1  Rohit Saxena  [^1] footnotemark: 1  Xiaotang Du [^1] footnotemark: 1  Ping Nie511footnotemark: 1  Yu Zhao [^1] footnotemark: 1 Laura Perez-Beltrachini1  Max Ryabinin4  Xuanli He3  Clémentine Fourrier2 Pasquale Minervini [^1] footnotemark: 1thanks: Corresponding authors. 1School of Informatics, University of Edinburgh  2 Hugging Face 3Department of Computer Science, University College London  4Together AI 5School of Electronics Engineering and Computer Science, Peking University {first.last, lperez, p.minervini}@ed.ac.uk  mryabinin0@gmail.com clementine@huggingface.co  xuanli.he@ucl.ac.uk   ping.nie@pku.edu.cn
@@ Acknowledgements
Experiments are being conducted mainly at the Edinburgh International Data Facility (EIDF) and on the internal clusters of the School of Informatics, University of Edinburgh. APG was supported by the United Kingdom Research and Innovation (grant EP/S02431X/1), UKRI Centre for Doctoral Training in Biomedical AI at the University of Edinburgh, School of Informatics. PM was partially funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1); an industry grant from Cisco; and a donation from Accenture LLP; and is grateful to NVIDIA for the GPU donations. XH and PM are funded by an industry grant from Cisco. GW was supported by ILCC program (School of Informatics Funding Package) at the University of Edinburgh, School of Informatics. RS, XD, and YZ are supported in part by the UKRI Centre for Doctoral Training in Natural Language Processing, funded by UK Research and Innovation (grant EP/S022481/1) and the School of Informatics.